OpenStudy (anonymous):

Limit as x approaches infinity of (16x/(4x+9))

OpenStudy (anonymous):

I.e \[\lim_{x \rightarrow \infty} \frac{ 16x }{ 4x+9 }\]

OpenStudy (dayakar):

divide numerator and denominator with x plz ,do it

OpenStudy (idealist10):

Apply L'Hopital's Rule. 16x/4x=4 Answer: 4

OpenStudy (dayakar):

r u there

OpenStudy (anonymous):

Ahh right

OpenStudy (anonymous):

Sorry, got disconnected But I see it now, both L'Hopitals Rule and Dayakar's method, (16x/(4x+9)) = 16/(4 + 9/x) --> 9/x approaches 0 as x approaches infinity --> the expression approaches 16/4 = 4. Thanks guys! Damn, really out of shape mathematically..